### v2

# Joint Channel, Data and Radar Parameter Estimation for AFDM Systems in Doubly-Dispersive Channels

Abstract
We propose new schemes for joint channel and data estimation (JCDE) and radar parameter estimation (RPE) in doubly-dispersive channels, such that integrated sensing and communications (ISAC) is enabled by user equipment (UE) independently performing JCDE, and base stations (BSs) performing RPE. The contributed JCDE and RPE schemes are designed for waveforms known to perform well in doubly-dispersive channels, under a unified model that captures the features of either legacy orthogonal frequency division multiplexing (OFDM), state-of-the-art (SotA) orthogonal time frequency space (OTFS), and next-generation affine frequency division multiplexing (AFDM) systems. The proposed JCDE algorithm is based on a Bayesian parametric bilinear Gaussian belief propagation (PBiGaBP) framework first proposed for OTFS and here shown to apply to all aforementioned waveforms, while the RPE scheme is based on a new probabilistic data association (PDA) approach incorporating a Bernoulli-Gaussian denoising, optimized via expectation maximization (EM). Simulation results demonstrate that JCDE in AFDM systems utilizing a single pilot per block significantly outperforms the SotA alternative even if the latter is granted a substantial power advantage. Similarly, the AFDM-based RPE scheme is found to outperform the OTFS-based approach, as well as the sparse Bayesian learning (SBL) technique, regardless of the waveform used.
Index Terms:
ISAC, AFDM, OTFS, JCDE, PBiGaBP, PDA, EM, RPE, Bayesian Inference.
I Introduction
The building blocks of communications and radar systems have been traditionally designed separately and assumed to utilize orthogonal resources. Integrated sensing and communications (ISAC)  [^1] Also commonly referred to as joint communication and sensing (JCAS) and joint communications and radar (JCR). [2, 3, 4], a recently-emerged technology seen by many as a key driver of future mobile communication systems, challenges that established concept by considering the simultaneous provision of radar and communications functionalities under the same architecture, having the advantages of more efficient use of resources, lower power consumption, decreased hardware costs and the potential for wide market penetration [5, 6, 7].
Staying true to the integrative aspect of the proposition [^2] We distinguish ISAC from earlier co-existence approaches [21, 22], without integration of these functions., earlier contributions on ISAC focused on schemes employing existing communications waveforms. In particular, many methods found in current literature seek to take advantage of certain features of channel estimation – especially in higher frequency bands where less scattering and higher resolution enable the estimation of individual channel paths – to incorporate radar-like functionalities, such as the estimation of radar parameters from echoes of the transmit signal [8, 9, 10, 11, 12].
Examples of the latter are techniques based on the orthogonal frequency division multiplexing (OFDM) waveform [10, 11] which, however, indicated that the high inter-carrier interference (ICI) can be a bottleneck that reduces robustness to high Doppler shifts. This factor alone, which stems from the fact that the channel diagonals spread into a decaying band [13], severely hinders the use of OFDM in high mobility scenarios [14], as envisioned for 6G systems.
Another potential candidate for the ISAC paradigm is the orthogonal time frequency space (OTFS) waveform, which was originally proposed in [15] as a novel two-dimensional (2D) modulation scheme that embeds information directly on the delay-Doppler (DD) domain. Since the DD domain had a direct parallel to common radar processing framework, OTFS was quite popular at its inception, attracting much work that demonstrated its effectiveness for ISAC [8, 9, 10, 12, 11]. Besides the significantly higher complexity compared to OFDM, which is inherent to its 2D modulation approach, another inherent limitation of OTFS is that it does not achieve the optimal diversity order in doubly-dispersive channels [16].
In response to the above, affine frequency division multiplexing (AFDM) has recently emerged as a strong candidate for ISAC [17, 18, 19, 20], gaining much attention thanks to its robust communications performance in high mobility scenarios and its ability to achieve optimal diversity order in doubly-dispersive channels. The key idea of AFDM is to employ the discrete affine Fourier transform (DAFT) to generate multiple orthogonal chirps that are then used to modulate the transmit signal. Since AFDM is otherwise similar to OFDM in structure, it does not suffer from the high modulation complexity of OTFS, and since the chirp parameters of AFDM can be optimized to match the maximum Doppler shift of the doubly-dispersive channel, a full DD representation of the channel can be achieved, giving the approach remarkable robustness to high mobility and/or frequency.
It has been in fact demonstrated that AFDM has a superior bit error rate (BER) performance compared to OFDM, achieving also BERs comparable to OTFS while requiring lower overhead for channel estimation [19]. In turn, evidence that AFDM can be effectively utilized for ISAC has also been given, for instance in [18] where a simple matched-filter (MF) approach was suggested. Building on the latter, an improved AFDM-base ISAC performance was also achieved in [20], albeit at the expense of a significant increase in computational complexity due to the exhaustive maximum likelihood (ML) grid search required.
One aspect of the ISAC concept that was often overlooked, and which recently has been addressed more frontally, however, is the impact of pilot overhead onto the overall approach. It has been shown in [19], for instance, that a major advantage of AFDM over OFDM and OTFS is the lighter requirement in the number of pilots, which in turn can be taken advantage of by enabling the base station (BS) to perform radar parameter estimation (RPE) without expensive self-interference cancellation (SIC) techniques. The trend to reduce pilot overhead suggests, therefore, the adoption of joint channel and data estimation (JCDE) technology, which has been shown to be very effective in allowing efficient channel estimation with very few pilots [^3] The embedded channel estimation method proposed in [19] requires a single pilot, but surrounded by a good amount of zeros, such that the fraction of the AFDM frame dedicated to payload is affected not just by the sole pilot but also the zero-padding. and under strenuous conditions [23, 24, 25, 26, 27, 1], significantly reduce the overhead of coherent wireless systems. One example of the latter is the Bayesian bilinear inference (BBI) JCDE scheme for OFDM systems proposed [27], which was shown to have the ability of handling channel aging and to entirely remove the need of pilots in favor of adequate coding of an equivalent rate. Another is the work in [1], where a Bayesian parametric bilinear Gaussian belief propagation (PBiGaBP) JCDE scheme for OTFS systems was proposed, which was shown to asymptotically approach the idealized scheme where perfect channel state information (CSI) knowledge is available.
In view of all the above, we propose in this paper a combined and generalized JCDE and ISAC solution for systems known to cope well with doubly-dispersive channels, including OFDM, OTFS and AFDM. The proposed solution comprises two main contributions. The first is a new JCDE, designed with basis on the Bayesian PBiGaBP originally proposed for OTFS waveforms in [1], which is shown to function also for the other modulation schemes under the general doubly dispersive channel model described in [13]. Comparisons of the state-of-the-art (SotA) OTFS and the contributed AFDM methods against a genie-aided linear Gaussian belief propagation (GaBP) benchmark scheme, in which perfect CSI is assumed, demonstrate that the proposed solution under AFDM achieves the same performance of the OTFS SotA, both of which also closely approach the idealistic benchmark, indicating that there is no penalty for the reduced modulation complexity advantage obtained when utilizing AFDM instead of OTFS, and that the proposed scheme is effective in reducing pilot overhead.
The second contribution is a new RPE scheme based on the vectorized probabilistic data association (PDA) approach [24, 28] incorporating a Bernoulli-Gaussian denoiser tuned via expectation maximization (EM) [29] which quickly converges to the appropriate distribution parameters.
Compared to the SotA alternative of [30], which is designed around the sparse Bayesian learning (SBL) algorithm, the contributed method is shown to outperform the latter over any of the relevant waveforms, namely, OFDM, OTFS in [30] and AFDM. The technique also has an inherent and substantial computational cost advantage over ML-based SotA alternatives such as those in [19] and [20].
The contributions of the article are summarized as follows:
  • In Section II, we concisely revise the unified OFDM, OTFS and AFDM system model for doubly-dispersive channels from [13].
  • Taking advantage of the latter, we propose in Section III a PBiGaBP-based scheme for JCDE, which applies to all the aforementioned waveforms.
  • We show that even if a single pilot per block is used, the proposed JCDE-AFDM system significantly outperforms both the OFDM and OTFS alternatives [1], as well as the AFDM method in [19].
  • Finally, in Section IV-B a novel PDA-based method for RPE in doubly-dispersive channels is proposed, which in addition to enabling monostatic implementation without requiring full-duplex (FD) technology, is shown to outperform the SotA of [30].
II System Model
We follow [20] and consider an ISAC scenario composed of a single-antenna [^4] The generalization of the proposed methods to the multiple-input multiple-output (MIMO) case requires the formulation and solution of a non-trivial transmit (TX) and receive (RX) beamforming optimization problem, with objective and constraints designed to satisfy both communication and sensing requirements, but which otherwise have no fundamental influence on the JCDE and RPE schemes that will be proposed. That contribution is therefore out of the scope of this article and will be offered in a follow-up work. BS transmitting in downlink and taking advantage of first-order reflected (echo) signals to perform RPE in order to detect the ranges and velocities of objects (targets) in its surroundings, while the user equipment (UE) performs JCDE as illustrated in Figure 1. As depicted in the figure, it is assumed that a total of $P_{\mathrm{B}}$ target-originated echo signals reach the BS, which might be either from UE or other objects in the ambient. In turn, each UE is assumed to receive the signal transmitted by the BS scattered onto $P_{\mathrm{U}}+1$ paths, with $p_{\mathrm{U}}=0$ denoting the line-of-sight (LoS), without loss of generality (wlg), and the remaining $P_{\mathrm{U}}\leq P_{\mathrm{B}}$ paths being non-LoS (NLoS) components of the multi-path channel.
For the sake of simplicity and again wlg, we shall hereafter attain to a single user, since the multi-user (MU) case has no impact onto the JCDE procedure performed independently at each user, and is transparent to the RPE algorithm, as long as the radar receiver has sufficient degrees-of-freedom (DoF) to resolve all $P_{\mathrm{B}}$ echoed signals it receives. Also for simplicity, although we differentiate $P_{\mathrm{B}}$ and $P_{\mathrm{U}}$ in Fig. 1, in the equations that follow, the number of paths/targets will be denoted by $P$, with the distinction between the two cases left to context.
$\mathbf{\Phi}_{p}\triangleq\text{diag}\bigg{(}[\overbrace{e^{-j2\ell_{p}\pi},e^{-j2(\ell_{p}-1)\pi},\dots,e^{-j4\pi},e^{-j2\pi}}^{\ell_{p}\;\text{terms}},\overbrace{1,1,\dots,1,1}^{N-\ell_{p}\;\text{ones}}]\bigg{)}\in\mathbb{C}^{N\times N}.\vspace{-2ex}$ (5)
$\bm{\Omega}\triangleq\text{diag}\bigg{(}[1,e^{-j2\pi/N},\dots,e^{-j2\pi(N-2)/N},e^{-j2\pi(N-1)/N}]\bigg{)}\in\mathbb{C}^{N\times N}.$ (6)
Finally, for the sake of enabling direct comparison with legacy (OFDM) and SotA (OTFS) approaches, it shall be assumed that the BS is capable of performing perfect SIC  [^5] The problem of SIC mitigation for the purpose of RPE in AFDM has already been addressed in [20], and solutions for the related problems of interference cancellation in orthogonal chirp division multiplexing (OCDM), or more generally for in-band interference cancellation in single antenna systems have also been proposed $e.g.$ in [31] and [32], respectively, which support the feasibility of the assumption. prior to RPE, to the advantage of the OFDM and OTFS approaches, since it has been shown [20] that SIC is more challenging and has higher computational complexity under those waveforms than under AFDM.
Figure 1: Illustration of an ISAC system with $P_{\mathrm{B}}$ target echoes received by the BS, and $P_{\mathrm{U}}$ resolvable paths seen at the UE.
II-A Generalized Doubly-Dispersive Channel Model Model
Consider a general doubly-dispersive wireless channel model [13] characterized by $1$ LoS and $P$ NLoS propagation paths, with each $p$-th path comprising an uncorrelated complex fading gain $h_{p}\in\mathbb{C}$, a path delay $\tau_{p}\in[0,\tau_{\text{max}}]$ and a Doppler shift $\nu_{p}\in[-\nu_{\text{max}},\nu_{\text{max}}]$, respectively. The delays and Doppler shifts of the doubly-dispersive channel are assumed to be bounded by a maximum delay $\tau_{\text{max}}$[s] and a maximum Doppler shift $\nu_{\text{max}}$[Hz], respectively, such that the time domain (TD) channel impulse response can be described by [33]
$h(t,\tau)\triangleq\sum_{p=0}^{P}h_{p}\cdot e^{j2\pi\nu_{p}t}\cdot\delta(\tau-\tau_{p}),$ (1)
where $t$ and $\tau$ are the continuous time and delay variables, respectively.
Leveraging equation (1), the input-output relationship of the channel in the TD can be obtained via the linear convolution of the transmit signal $s(t)$ and channel impulse response as described in [33, eq. (10.69)], which added with noise yields
$r(t)=s(t)*h(t,\tau)+w(t)=$
[…]
where $r(t)$ and $w(t)$ are the received signal and additive white Gaussian noise (AWGN), respectively.
Defining $r[n]$ and $s[n]$, with $n\in\{0,\dots,N-1\}$, to be the sampled sequences of $r(t)$ and $s(t)$, respectively, with samples taken at the rate $f_{\mathrm{S}}\triangleq\frac{1}{T_{\mathrm{S}}}$[Hz], the following discrete equivalent of equation (II-A) can be obtained
$r[n]=\sum_{\ell=0}^{\infty}s[n-\ell]\bigg{(}\sum_{p=0}^{P}h_{p}\cdot e^{j2\pi\frac{\nu_{p}}{f_{\mathrm{S}}}n}\cdot\delta\Big{(}\ell-\frac{\tau_{p}}{T_{\mathrm{S}}}\Big{)}\bigg{)}+w[n],$ (3)
where $\ell$ is the discrete delay index, and $T_{\mathrm{S}}$ the delay resolution as well as the sampling interval, such that $\tau=\ell\cdot T_{\mathrm{S}}$.
Finally, defining the normalized digital Doppler shift and the normalized delay of the $p$-th path respectively as $f_{p}\triangleq\frac{N\nu_{p}}{f_{\mathrm{S}}}$ and $\ell_{p}\triangleq\frac{\tau_{p}}{T_{\mathrm{S}}}$, where the delay resolution $T_{\mathrm{S}}$ is assumed to be sufficiently high to ensure that $\ell_{p}-\lfloor\frac{\tau_{p}}{T_{\mathrm{S}}}\rceil\approx 0$, the discrete convolution described in equation (3) can be rewritten in terms of the circular convolution
$\mathbf{r}=\bigg{(}\sum_{p=0}^{P}h_{p}\cdot\mathbf{\Phi}_{p}^{\phi_{\text{CP}}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}}\bigg{)}\cdot\mathbf{s}+\mathbf{w}=\mathbf{H}\cdot\mathbf{s}+\mathbf{w},\vspace{-0.5ex}$ (4)
where the inclusion and posterior removal of a cyclic prefix (CP) of length $N_{\text{CP}}$ to the sequence $s[n]$ is already taken into account [13].
In equation (4), $\mathbf{r}\triangleq[r[0],\cdots,r[N-1]]\in\mathbb{C}^{N\times 1}$ and $\mathbf{s}\triangleq[s[0],\cdots,s[N-1]]\in\mathbb{C}^{N\times 1}$ are the transmit and received signal sample vectors; $\mathbf{w}\in\mathbb{C}^{N\times 1}$ denotes AWGN; $\mathbf{H}\in\mathbb{C}^{N\times N}$ is the effective channel matrix; $\mathbf{\Phi}_{p}\in\mathbb{C}^{N\times N}$, described in equation (5), is such that $\mathbf{\Phi}_{p}^{{}_{\text{CP}}}$ captures the result of the inclusion of the CP onto the $p$-th channel path; $\bm{\Omega}\in\mathbb{C}^{N\times N}$, described in equation (6), is a diagonal matrix containing the $N$ complex roots of one and $\mathbf{\Pi}\in\{0,1\}^{N\times N}$ is the forward cyclic shift matrix [^6] To clarify, the matrix $\mathbf{\Pi}$ is such that $\bm{A}\cdot\mathbf{\Pi}^{\ell_{p}}$, with $\ell_{p}\in\mathbb{N}_{0}$, is a cyclic left-shifted version of the matrix $\bm{A}$, such that the first $\ell_{p}$ columns of $\bm{A}$ are moved to the positions of the last $\ell_{p}$ columns. Notice also that $\mathbf{\Pi}^{0}$ is the $N\times N$ identity matrix, such that $\bm{A}\cdot\mathbf{\Pi}^{0}=\bm{A}$., with elements given by
$\pi_{i,j}=\delta_{i,j+1}+\delta_{i,j-(N-1)}\;\;\text{where}\;\;\delta_{ij}\triangleq\begin{cases}0&\text{if }i\neq j,\\ 1&\text{if }i=j.\end{cases}\vspace{-0.25ex}$ (7)
Notice that the channel $\mathbf{H}$ implicitly defined in (4) is a general representation of a doubly-dispersive channel [33], which shall be exploited in the sequel to formulate estimation and detection models over various waveforms.
II-B OFDM Signal Model
Let $\mathcal{C}\in\mathbb{C}^{Q\times 1}$ denote an arbitrary complex constellation of cardinality $Q$ and average energy $E_{\mathrm{S}}$, associated with a given digital modulation scheme. In OFDM, an input information vector $\mathbf{x}\in\mathcal{C}^{N\times 1}$ containing $N$ symbols taken from $\mathcal{C}$ are modulated into the transmit signal
$\mathbf{s}_{\text{OFDM}}=\mathbf{F}_{N}^{\mathsf{H}}\cdot\mathbf{x}\in\mathbb{C}^{N\times 1},$ (8)
where $\mathbf{F}_{N}$ denotes the $N$-point normalized discrete Fourier transform (DFT) matrix.
$\bm{\varPhi}_{p}\triangleq\text{diag}\bigg{(}[\overbrace{e^{-j2\pi(N^{2}-2N\ell_{p})},e^{-j2\pi(N^{2}-2N(\ell_{p}-1))},\dots,e^{-j2\pi(N^{2}-2N)}}^{\ell_{p}\;\text{terms}},\overbrace{1,1,\dots,1,1}^{N-\ell_{p}\;\text{ones}}]\bigg{)}\in\mathbb{C}^{N\times N}.\vspace{-1ex}$ (22)
After undergoing the circular convolution with the doubly-dispersive channel, the corresponding OFDM receive signal can be written in a form similar to that of equation (4), namely
$\mathbf{r}_{\text{OFDM}}\triangleq\mathbf{H}\cdot\mathbf{s}_{\text{OFDM}}+\mathbf{w}\in\mathbb{C}^{N\times 1},\vspace{-0.5ex}$ (9)
which can be demodulated into
$\mathbf{y}_{\text{OFDM}}=\mathbf{F}_{N}\cdot\mathbf{r}_{\text{OFDM}}\in\mathbb{C}^{N\times 1}.\vspace{-0.5ex}$ (10)
From equations (9) and (10), one obtains a straightforward input-output relationship for OFDM modulation, namely
$\mathbf{y}_{\text{OFDM}}=\mathbf{G}_{\text{OFDM}}\cdot\mathbf{x}+\tilde{\mathbf{w}}\in\mathbb{C}^{N\times 1},\vspace{-0.5ex}$ (11)
where $\tilde{\mathbf{w}}\triangleq\mathbf{F}_{N}\mathbf{w}\in\mathbb{C}^{N\times 1}$ is an equivalent AWGN vector with the same statistics as $\mathbf{w}$, while $\mathbf{G}_{\text{OFDM}}$ is the effective OFDM channel, defined by
$\mathbf{G}_{\text{OFDM}}\triangleq\sum_{p=0}^{P}h_{p}\cdot\mathbf{F}_{N}\cdot(\mathbf{\Phi}_{p}^{\phi_{\text{CP}}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot\mathbf{F}_{N}^{\mathsf{H}}.$ (12)
II-C OTFS Signal Model
Consider an OTFS system in which an input information matrix $\mathbf{X}\in\mathcal{C}^{K\times M}$, containing $K\times M$ complex symbols taken from an arbitrary constellation $\mathcal{C}$, is modulated into the transmit signal [^7] For simplicity, we assume that all pulse-shaping operations are perfect ($i.e.,$ the corresponding pulse-shaping waveforms are rectangular), such that the corresponding sample matrices can be reduced to identity matrices.
$\mathbf{s}_{\text{OTFS}}\triangleq\text{vec}(\mathbf{S})=(\mathbf{F}_{M}^{\mathsf{H}}\otimes\mathbf{I}_{K})\cdot\mathbf{x}\in\mathbb{C}^{KM\times 1},$ (13)
where $\text{vec}(\cdot)$ denotes the vectorization of a matrix, obtained by stacking the columns; the symbol $\otimes$ denotes Kronecker product; and $\mathbf{S}$ is a TD symbol matrix obtained [^8] Equivalently, $\mathbf{S}$ can be obtained as the Heisenberg transform of the inverse symplectic finite Fourier transform (ISFFT) of $\mathbf{X}$, $i.e.$ $\mathbf{S}=\mathbf{F}_{K}^{\mathsf{H}}\mathbf{X}_{\text{FT}}$ with $\mathbf{X}_{\text{FT}}\triangleq\mathbf{F}_{K}\mathbf{X}\mathbf{F}_{M}^{\mathsf{H}}\in\mathbb{C}^{K\times M}$. from the inverse discrete Zak transform (IDZT) of $\mathbf{X}$, $i.e.$ [15]
$\mathbf{S}=\mathbf{X}\mathbf{F}_{M}^{\mathsf{H}}\in\mathbb{C}^{K\times M}.$ (14)
We highlight that the notation in equation (13) is in line with the strategy described $e.g.$ in [34], whereby the OTFS signals are first vectorized and then appended with a CP of length $N_{CP}$ in order to eliminate inter-frame interference, in similarity with OFDM. Taking advantage of such similarity, and in order to allow for direct comparisons between the two waveforms, we shall hereafter set $K\times M=N$.
After transmission over the doubly-dispersive channel $\mathbf{H}$, the OTFS receive signal can be modeled as
$\mathbf{r}_{\text{OTFS}}\triangleq\mathbf{H}\cdot\mathbf{s}_{\text{OTFS}}+\mathbf{w}\in\mathbb{C}^{N\times 1},$ (15)
which is analogous to the OFDM model of equation (11).
Unlike OFDM, however, the detection of the information symbols in $\mathbf{x}$ from $\mathbf{r}_{\text{OTFS}}$ requires reversing the vectorization and the IDZT operations employed in the construction of $\mathbf{s}_{\text{OTFS}}$, resulting in a distinct effective channel.
In particular, let $\mathbf{R}\triangleq\text{vec}^{-1}(\mathbf{r}_{\text{OTFS}})\in\mathbb{C}^{K\times M}$, where $\text{vec}^{-1}(\cdot)$ denotes the de-vectorization operation whereby a vector of size $KM\times 1$ is reshaped into a matrix of size $K\times M$, and consider the corresponding discrete Zak transform (DZT)  [^9] Equivalently, $\mathbf{Y}$ can be obtained as the symplectic finite Fourier transform (SFFT) of the Wigner transform $\mathbf{Y}_{\text{FT}}\triangleq\mathbf{F}_{K}\mathbf{R}$ of the matrix $\mathbf{R}$, which yields $\mathbf{Y}=\mathbf{F}_{K}^{\mathsf{H}}\mathbf{Y}_{\text{FT}}\mathbf{F}_{M}\in\mathbb{C}^{K\times M}$.
$\mathbf{Y}=\mathbf{R}\mathbf{F}_{M}\in\mathbb{C}^{K\times M}.$ (16)
The corresponding demodulated OTFS signal is given by
$\mathbf{y}_{\text{OTFS}}\triangleq\text{vec}(\mathbf{Y})=(\mathbf{F}_{M}\otimes\mathbf{I}_{K})\cdot\mathbf{r}_{\text{OTFS}}\in\mathbb{C}^{N\times 1},$ (17)
which in turn can be rewritten in the form of the direct input-output relation
$\mathbf{y}_{\text{OTFS}}=\mathbf{G}_{\text{OTFS}}\cdot\mathbf{x}+\tilde{\mathbf{w}}_{\text{OTFS}}\in\mathbb{C}^{N\times 1},$ (18)
where $\tilde{\mathbf{w}}_{\text{OTFS}}\triangleq(\mathbf{F}_{M}\otimes\mathbf{I}_{K})\cdot\mathbf{w}\in\mathbb{C}^{N\times 1}$ is an equivalent AWGN vector with the same statistics as $\mathbf{w}$, while $\mathbf{G}_{\text{OTFS}}\in\mathbb{C}^{N\times N}$ is the effective OTFS channel in the DD domain, defined by
$\mathbf{G}_{\text{OTFS}}\triangleq\sum_{p=0}^{P}h_{p}\cdot(\mathbf{F}_{M}\otimes\mathbf{I}_{K})\cdot(\mathbf{\Phi}_{p}^{\phi_{\text{CP}}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot(\mathbf{F}_{M}^{\mathsf{H}}\otimes\mathbf{I}_{K}).$ (19)
Comparing equations (12) and (19), one can appreciate how the channel modeling approach of [13] elucidates both the similarity in form, and the distinction in effect between the OFDM and OTFS waveforms in doubly-dispersive channels.
II-D AFDM Signal Model
Finally, let us describe the input-output relationship of the AFDM waveform associated with a doubly-dispersive channel. Similarly to the above, letting $\mathbf{x}$ denote the information vector with symbols taken from the constellation $\mathcal{C}$, the corresponding AFDM modulated transmit signal is given by its inverse discrete affine Fourier transform (IDAFT), $i.e.$
$\mathbf{s}_{\text{AFDM}}=(\mathbf{\Lambda}_{1}^{\mathsf{H}}\mathbf{F}_{N}^{\mathsf{H}}\mathbf{\Lambda}_{2}^{\mathsf{H}})\cdot\mathbf{x}\in\mathbb{C}^{N\times 1},$ (20)
with the matrices $\mathbf{\Lambda}_{i}$ defined as
$\mathbf{\Lambda}_{i}\triangleq\text{diag}\big{(}\big{[}1,\cdots,e^{-j2\pi c_{i}n^{2}},\cdots,e^{-j2\pi c_{i}(N-1)^{2}}\big{]}\big{)}\in\mathbb{C}^{N\times N},$ (21)
where the constants $c_{1}$ and $c_{2}$ are chosen appropriately to match the maximum Doppler of the channel, and include the insertion of a chirp-periodic prefix (CPP) to mitigate the effects of multipath propagation [19].
It was shown in [13] that after going through a doubly-dispersive channel, an AFDM modulated symbol vector as given in equation (20) with the inclusion of a CPP as described in [19], can be modeled similarly to equation (4), only with the CP matrix $\bm{\Phi}_{p}$ of equation (5) replaced by the corresponding CPP matrix $\bm{\varPhi}_{p}$ as given in equation (22), such that the corresponding received signal is given by
$\mathbf{r}_{\text{AFDM}}\triangleq\mathbf{H}\cdot\mathbf{s}_{\text{AFDM}}+\mathbf{w}\in\mathbb{C}^{N\times 1}.$ (23)
Demodulating the signal in equation (23) yields
$\mathbf{y}_{\text{AFDM}}=(\mathbf{\Lambda}_{2}\mathbf{F}_{N}\mathbf{\Lambda}_{1})\cdot\mathbf{r}_{\text{AFDM}}\in\mathbb{C}^{N\times 1},$ (24)
which in turn can be expressed as
$\mathbf{y}_{\text{AFDM}}=\mathbf{G}_{\text{AFDM}}\cdot\mathbf{x}+\tilde{\mathbf{w}}_{\text{AFDM}}\in\mathbb{C}^{N\times 1},$ (25)
where $\tilde{\mathbf{w}}_{\text{AFDM}}\triangleq(\mathbf{\Lambda}_{2}\mathbf{F}_{N}\mathbf{\Lambda}_{1})\cdot\mathbf{w}\in\mathbb{C}^{N\times 1}$ is an equivalent AWGN vector with the same statistics as $\mathbf{w}$, and $\mathbf{G}_{\text{AFDM}}\in\mathbb{C}^{N\times N}$ in the effective AFDM channel matrix, given by
$\mathbf{G}_{\text{AFDM}}\triangleq\sum_{p=0}^{P}h_{p}\cdot(\mathbf{\Lambda}_{2}\mathbf{F}_{N}\mathbf{\Lambda}_{1})\cdot(\bm{\varPhi}_{p}^{c_{1}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}})\cdot(\mathbf{\Lambda}_{1}^{\mathsf{H}}\mathbf{F}_{N}^{\mathsf{H}}\mathbf{\Lambda}_{2}^{\mathsf{H}}),$ (26)
where we emphasize that $\bm{\varPhi}_{p}$ is as in equation (22).
Once again, it is clear that equations (12), (19) and (26) have the same structure, as do the input-output relationships described by equations (11), (18) and (25), such that the JCDE and RPE techniques proposed in the sequel are applicable to OFDM, OTFS and AFDM waveforms alike.
III Joint Channel and Data Estimation
In this section, we introduce the proposed JCDE method for OFDM, OTFS and AFDM systems, under the general system model described in the previous section, and compare their relative performances. For the sake of convenience, we shall express the effective channels for the three waveforms given in equations (12), (19) and (26) in the form
$\mathbf{y}=\sum_{p=0}^{P}h_{p}\cdot\mathbf{\Gamma}_{p}\cdot\mathbf{x}+\tilde{\mathbf{w}}\in\mathbb{C}^{N\times 1},$ (27)
where the matrices $\mathbf{\Gamma}_{p}$ captures the long-term delay-Doppler statistics of the channel, which are assumed to be known.
From equation (27), it is evident that each received signal in $\mathbf{y}$ can described by
$y_{n}=\sum_{p=0}^{P}\sum_{m=0}^{N-1}h_{p}\cdot\gamma_{p:n,m}\cdot x_{m}+\tilde{w}_{n},$ (28)
where we slightly deviate from the notation employed earlier in equation (3) by denoting the $m$-th information symbol as $x_{m}$ instead of $x[m]$, and accordingly the $n$-th receive signal sample by $y_{n}$ instead of $y[n]$, for future convenience.
III-A JCDE via Parametric Bilinear Gaussian Belief Propagation
We have first proposed the PBiGaBP framework for JCDE under OTFS signaling in [1]. Thanks to the generalized channel model described above, however, it is obvious that the same technique applies to OFDM and AFDM as well. We therefore proceed to succinctly describe the approach, referring the reader to [1] for further details.
For starters, referring to Figure 2, it is initially assumed [^10] Later we shall adopt the strategy proposed in [19] and consider, for the AFDM case, the alternative of utilizing only one pilot followed by $B-1$ zeros among the $N$ symbols in $\mathbf{x}$, for the purpose of enabling the RPE scheme of Section IV-B. that in any of the three systems compared, a block of $B$ pilots [^11] We use the first column of the $B\times B$ Zadoff-chu sequence [1] as our pilot column. is embedded in the $N\times 1$ symbol vector $\mathbf{x}$.
We shall also clarify that although the placement of pilots amidst payload (data) symbols can be optimized under various criteria [35], in the context of our discussion where issues such as distortions due hardware imperfection [36], pilot contamination [37], and channel aging [27] are not considered, the location of pilots has no effect on the performance of the JCDE. Suffice it therefore, to consider the amount of power allocated to pilots, relative to that allocated to data symbols.
With these issues addressed, the PBiGaBP framework for JCDE introduced here can be summarized as follows:
  • Soft interference cancellation (soft IC), which consists of removing the inter-symbol interference from $\mathbf{y}$, by using temporary estimates (soft replicas) of $\mathbf{h}$ and $\mathbf{x}$ generated in the previous iteration;
  • Belief generation (BG), whereby the likelihood (beliefs) of the estimates of $\mathbf{h}$ and $\mathbf{x}$ are obtained from $\mathbf{y}$ after soft IC and under scalar Gaussian approximation (SGA);
  • Soft replica generation (soft RG), whereby soft replicas are generated from the beliefs via conditional expectations.
The signal processing operations corresponding to each of these steps are described below.
III-A1 Soft IC
Let us consider the $i$-th iteration of the algorithm, and denote the soft replicas of the $m$-th symbol and the $p$-th channel gain associated with the $n$-th receive signal $y_{n}$, at the previous iteration respectively by $\hat{x}_{n,m}^{(i-1)}$ and $\hat{h}_{n,p}^{(i-1)}$. Then, the mean-squared-errors (MSEs) of these estimates computed for the $i$-th iteration are given by
$\hat{\sigma}^{2(i)}_{x:{n,m}}\triangleq\mathbb{E}_{x}\big{[}|x-\hat{x}_{n,m}^{(i-1)}|^{2}\big{]}=E_{\mathrm{S}}-|\hat{x}_{n,m}^{(i-1)}|^{2},\forall(n,m),$ (29a)
$\hat{\sigma}^{2(i)}_{h:{n,p}}\triangleq\mathbb{E}_{h}\big{[}|h-\hat{h}_{n,p}^{(i-1)}|^{2}\big{]}=\sigma_{h}^{2}-|\hat{h}_{n,p}^{(i-1)}|^{2},\forall(n,p),$ (29b)
where $\mathbb{E}_{x}$ refers to expectation over the all possible symbols $x$ in the constellation $\mathcal{C}$, while $\mathbb{E}_{h}$ refers to expectation over all possible outcomes of $h\sim\mathcal{CN}(0,\sigma^{2}_{h})$, respectively [^12] Although the average power of each channel path may be generally different, especially in sub-6GHz channels where the propagation lengths of different paths can vary from meters to kilometers, it is commonly assumed in high-frequency (mmWave and TeraHz) channels, where ranges are much shorter, that the gains of all paths follow the same distribution [18, 19, 20]..
At a given $i$-th iteration, the objective of the soft IC step is to compute the symbol- and channel-centric replicas $\tilde{y}_{x:{m,n}}^{(i)}$ and $\tilde{y}_{h:{p,n}}^{(i)}$ and the corresponding variances $\tilde{\sigma}^{2(i)}_{x:{m,n}}$ and $\tilde{\sigma}^{2(i)}_{h:{p,n}}$, using the soft estimates $\hat{x}_{n,m}^{(i-1)}$ and $\hat{h}_{n,p}^{(i-1)}$ and their variances $\hat{\sigma}_{x:{n,m}}^{2(i-1)}$ and $\hat{\sigma}^{2(i-1)}_{h:{n,p}}$ obtained in the previous iteration.
In view of equation (28), we straightforwardly have
$\tilde{y}_{x:{m,n}}^{(i)}=y_{n}-\sum_{p=0}^{P}\sum_{q\neq m}^{N-1}\hat{h}_{n,p}^{(i-1)}\cdot\gamma_{p:n,q}\cdot\hat{x}_{n,q}^{(i-1)}.$ (30)
These soft IC receive signal replicas follow Gaussian probability density functions (PDFs), namely
$\tilde{y}_{x:{m,n}}^{(i)}\sim p_{\text{y}_{n}|x_{m}}(\text{y}_{n}|x_{m})\propto\text{exp}\bigg{(}\!-\frac{|\text{y}_{n}\!-\!\tilde{\gamma}_{x:{m,n}}^{(i)}x_{m}|^{2}}{\tilde{\sigma}^{2(i)}_{x:{m,n}}}\bigg{)},$ (31)
where $\text{y}_{n}$ is an auxiliary variable and $\tilde{\gamma}_{x:{m,n}}^{(i)}$ are the corresponding soft IC effective channel gains, defined as
$\tilde{\gamma}_{x:{m,n}}^{(i)}\triangleq\sum_{p=0}^{P}\hat{h}_{n,p}^{(i-1)}\gamma_{p:n,m},$ (32)
while $\tilde{\sigma}^{2(i)}_{x:{m,n}}$ are the soft IC conditional variances given by
$\tilde{\sigma}^{2(i)}_{x:{m,n}}\triangleq\mathbb{E}_{x\neq x_{m},h}\big{[}|\tilde{y}_{x:{m,n}}^{(i)}-\tilde{\gamma}_{x:{m,n}}^{(i)}x_{m}|^{2}\big{]}$
$\approx\sum_{p=0}^{P}\hat{\sigma}^{2(i-1)}_{h:{n,p}}|\hat{y}^{(i)}_{x_{m}:n,p}|^{2}+\sum_{q\neq m}^{N-1}\hat{\sigma}_{x:{n,q}}^{2(i-1)}|\tilde{\gamma}_{x:q,n}^{(i)}|^{2}+N_{0}+$
$\sum_{p=0}^{P}\hat{\sigma}^{2(i-1)}_{h:{n,p}}\sum_{q\neq m}^{N-1}\hat{\sigma}_{x:{n,q}}^{2(i-1)}|\gamma_{p:n,q}|^{2}\!+\!E_{\mathrm{S}}\sum_{p=0}^{P}\hat{\sigma}^{2(i-1)}_{h:{n,p}}|\gamma_{p:n,m}|^{2},$
with $\hat{y}^{(i)}_{x_{m}:n,p}$ denoting the received signal estimate after cancellation of the $m$-th soft symbol estimate, which is given by
$\hat{y}^{(i)}_{x_{m}:n,p}\triangleq\sum_{q\neq m}^{N-1}\gamma_{p:n,q}\hat{x}_{n,q}^{(i-1)}.$ (34)
Similarly, the channel-centric replica can be computed as
$\tilde{y}_{h:{p,n}}^{(i)}=y_{n}-\sum_{q\neq p}^{P}\hat{h}_{n,q}^{(i)}\cdot\hat{y}_{h:{n,q}}^{(i)},$ (35)
with the corresponding variance given by
$\tilde{\sigma}^{2(i)}_{h:{p,n}}\!\triangleq\!\sum_{q\neq p}^{P}\!\hat{\sigma}_{h:{n,q}}^{2(i-1)}|\hat{y}_{h:{n,q}}^{(i)}|^{2}\!+\!\!\!\sum_{m=0}^{N-1}\!\!\hat{\sigma}^{2(i-1)}_{x:{n,m}}|\tilde{\gamma}_{h_{p}:{m,n}}^{(i)}|^{2}\!+\!N_{0}+\!\!$
$\sum_{q\neq p}^{P}\!\!\hat{\sigma}_{h:{n,q}}^{2(i)}\!\!\sum_{m=0}^{N-1}\!\!\hat{\sigma}^{2(i)}_{x:{n,m}}|\gamma_{q,nm}|^{2}\!+\!\sigma^{2}_{h}\sum_{m=0}^{N-1}\hat{\sigma}^{2(i)}_{x:{n,m}}|\gamma_{p:n,m}|^{2}\!,$
where $\hat{y}_{h:{n,p}}^{(i)}$ is the channel-centric soft channel estimate and $\tilde{\gamma}_{h_{p}:{m,n}}^{(i)}$ is the corresponding soft IC effective channel gain of the $p$-th path, respectively given by
$\hat{y}_{h:{n,p}}^{(i)}\triangleq\sum_{m=0}^{N-1}\gamma_{p:n,m}\hat{x}_{n,m}^{(i-1)}\;\,\text{and}\;\,\tilde{\gamma}_{h_{p}:{m,n}}^{(i)}\triangleq\sum_{q\neq p}^{P}\hat{h}_{n,q}^{(i-1)}\gamma_{q,nm}.$ (37)
III-A2 Belief Generation
In order to generate the beliefs of all symbols, consider the effect of substituting equation (28) into (30) such that the SGA holds under the assumption that $N\times P$ is a sufficiently large number, and that the individual estimation errors in $\hat{h}_{n,p}^{(i-1)}$ and $\hat{x}_{n,q}^{(i-1)}$ are independent.
Then, the belief corresponding to the $m$-th symbol $x_{m}$ at the $n$-th factor node is obtained by combining the contributions of all signals in $\mathbf{y}$, excluding $y_{n}$, under the PDF
$\!\!\!\!p_{\text{x}|x_{m}}(\text{x}|x_{m})\!=\!\prod_{q\neq n}^{N}\!p_{\text{y}_{q}|\text{x}_{m}}(\text{y}_{q}|x_{m})\propto\text{exp}\bigg{(}\!\!\!-\!\frac{|\text{x}\!-\!\tilde{x}_{n,m}^{(i)}|^{2}}{\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}}\!\bigg{)},\!$ (38a)
such that the desired beliefs and their variances are given by
$\tilde{x}_{n,m}^{(i)}\!\!\triangleq\!\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\!\sum_{q\neq n}^{N}\!\!\frac{\tilde{\gamma}_{x:{m,q}}^{*(i)}\tilde{y}_{x:{m,q}}^{(i)}}{\tilde{\sigma}_{x:{m,q}}^{2(i)}}\;\,\text{and}\;\,\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\!\!\triangleq\!\!\bigg{(}\!\sum_{q\neq n}^{N}\!\!\frac{|\tilde{\gamma}_{x:{m,q}}^{(i)}|^{2}}{\tilde{\sigma}_{x:{m,q}}^{2(i)}}\bigg{)}^{\!\!\!-1}\!\!\!\!.$ (39)
Similarly, the extrinsic beliefs of the channel gains can be shown to follow the approximate distribution
$p_{\text{h}|h_{p}}(\text{h}|h_{p})\propto\text{exp}\bigg{(}-\frac{|\text{h}-\!\tilde{\;\mathrm{h}}_{n,p}^{(i)}|^{2}}{\tilde{\sigma}_{\!\tilde{\;\mathrm{h}}:{n,p}}^{2(i)}}\bigg{)},$ (40)
with corresponding beliefs and variances given by
Figure 2: Illustration of pilot allocation schemes in OFDM, OTFS and AFDM.
$\!\tilde{\;\mathrm{h}}_{n,p}^{(i)}\!\triangleq\!\tilde{\sigma}_{\!\tilde{\;\mathrm{h}}:{n,p}}^{2(i)}\sum_{q\neq n}^{N}\frac{\hat{y}_{h:{qp}}^{*(i)}\tilde{y}_{h:{p,q}}^{(i)}}{\tilde{\sigma}_{h:{p,q}}^{2(i)}}\;\,\text{and}\;\,\tilde{\sigma}_{\!\tilde{\;\mathrm{h}}:{n,p}}^{2(i)}\!\triangleq\!\!\bigg{(}\!\sum_{q\neq n}^{N}\frac{|\hat{y}_{h:{qp}}^{(i)}|^{2}}{\tilde{\sigma}_{h:{p,q}}^{2(i)}}\bigg{)}^{\!\!-1}\!\!\!\!\!.$ (41)
III-A3 Soft RG
Finally, the soft replicas of $x_{m}$ and $h_{p}$ can be obtained from the conditional expectation given the extrinsic beliefs, again thanks to the SGA and under the assumption that the effective noise components in $\hat{x}_{n,m}^{(i)},\forall(n,m)$ and $\hat{h}_{n,p}^{(i)},\forall(n,p)$ are uncorrelated. In particular, we have
$\hat{x}_{n,m}^{(i)}=\frac{\sum\limits_{x\in\mathcal{C}}x\cdot p_{\text{x}|x}\big{(}\text{x}|x;\tilde{x}_{n,m}^{(i)},\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\big{)}\cdot p_{x}(x)}{\sum\limits_{x^{\prime}\in\mathcal{C}}p_{\text{x}|x^{\prime}}\big{(}\text{x}|x^{\prime};\tilde{x}_{n,m}^{(i)},\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\big{)}\cdot p_{x^{\prime}}(x^{\prime})},$ (42)
$\hat{h}_{n,p}^{(i)}=\frac{\int h\cdot p_{\text{h}|h}\big{(}\text{h}|h;\tilde{h}_{n,p}^{(i)},\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}\big{)}\cdot p_{h}(h)}{\int p_{\text{h}|h^{\prime}}\big{(}\text{h}|h^{\prime};\tilde{h}_{n,p}^{(i)},\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}\big{)}\cdot p_{h^{\prime}}(h^{\prime})},$ (43)
where $p_{\text{x}|x}\big{(}\text{x}|x;\tilde{x}_{n,m}^{(i)},\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}\big{)}$ and $p_{\text{h}|h}\big{(}\text{h}|h;\tilde{h}_{n,p}^{(i)},\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}\big{)}$ are the likelihood functions of the data and channel beliefs given in equations (38a) and (40), respectively, only with the notation extended to explicitly show the parameters (means and variances) computed above.
In order to compute the beliefs $\hat{x}_{n,m}^{(i)}$ with basis on equation (42), suffice it to notice that under the assumption that each element $x$ of the data vector is independently drawn with equal probability out of the $Q$-points of the constellation $\mathcal{C}$, the likelihood $p_{x}(x)$ is a Multinomial distribution of the $Q$-th order, such that for quadrature phase-shift keying (QPSK) modulation, we have
$\hat{x}_{n,m}^{(i)}\!=\!c_{x}\!\cdot\!\bigg{(}\!\text{tanh}\!\bigg{[}2c_{x}\frac{\Re\{{\tilde{x}_{n,m}^{(i)}}\}}{\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}}\bigg{]}\!\!+\!j\text{tanh}\!\bigg{[}2c_{x}\frac{\Im\{{\tilde{x}_{n,m}^{(i)}}\}}{\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}}\bigg{]}\!\bigg{)},\!\!$ (44)
where $c_{x}\triangleq\sqrt{E_{\mathrm{S}}/2}$ represents the magnitude of the real and imaginary components of the QPSK symbols.
After obtaining $\hat{x}_{n,m}^{(i)}$ as per equation (44), the final output is computed by damping the result in order to improve convergence [38], yielding
$\hat{x}_{n,m}^{(i)}=\beta_{x}\hat{x}_{n,m}^{(i)}+(1-\beta_{x})\hat{x}_{n,m}^{(i-1)}.$ (45)
In turn, the corresponding variance $\hat{\sigma}^{2(i)}_{x:{n,m}}$ is first updated following equation (29a), but with the updated symbol estimate, $i.e.$, $\hat{\sigma}^{2(i)}_{x:{n,m}}=E_{\mathrm{S}}-|\hat{x}_{n,m}^{(i)}|^{2}$, and then damped via
$\hat{\sigma}^{2(i)}_{x:{n,m}}=\beta_{x}\hat{\sigma}_{x:{n,m}}^{2(i)}+(1-\beta_{x})\hat{\sigma}_{x:{n,m}}^{2(i-1)}.$ (46)
Algorithm 1 Proposed PBiGaBP-based JCDE Method
Input: receive signal vector $\mathbf{y}$, pilot symbol vector $\mathbf{x}_{p}$, delay-Doppler matrices $\mathbf{\Gamma}_{p}$, maximum number of iterations $i_{\text{max}}$, total constellation energy $E_{\mathrm{S}}$, noise power $N_{0}$, average channel power per path $\sigma^{2}_{h}$, and damping factors $\beta_{x}$ and $\beta_{h}$.
Output: decoded symbols $\mathbf{x}_{d}$ and channel estimates $\hat{h}_{p},\forall p$.
1:
2:Initialization
3: - Set iteration counter to $i=0$ and amplitudes $c_{x}=\sqrt{E_{\mathrm{S}}/2}$
4: - Fix pilots to $\hat{x}_{n,m}^{(i)}=[\mathbf{x}_{p}]_{m}$ and set corresponding variances
5: to $\hat{\sigma}^{2(i)}_{x:{n,m}}=0,\forall n,m\in\mathcal{M}_{p}$
6: - Set initial data estimates to $\hat{x}_{n,m}^{(0)}=0$ and corresponding
7: variances to $\hat{\sigma}^{2(0)}_{x:{n,m}}=E_{\mathrm{S}},\forall n,m\in\mathcal{M}_{d}$
8: - Set initial channel estimates to $\hat{h}_{n,p}^{(0)}=0$ and corresponding
9: variances to $\hat{\sigma}^{2(0)}_{h:{n,p}}=\sigma_{h}^{2},\forall n,p$
10:
11:for $i=1$ to $i_{\text{max}}$ do
12:Channel Estimation: $\forall n,p$
13:Compute the variables $\hat{y}_{h:{n,p}}^{(i)}$ and $\tilde{\gamma}_{h_{p}:{m,n}}^{(i)}$ from eq. (37).
14:Compute soft signal $\tilde{y}_{h:{p,n}}^{(i)}$ from equation (35).
15:Compute soft signal variance $\tilde{\sigma}^{2(i)}_{h:{p,n}}$ from equation (III-A1).
16:Compute extrinsic channel belief $\tilde{h}_{n,p}^{(i)}$ and its variance $\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}$ from equation (41).
17:Compute denoised and damped channel estimate $\hat{h}_{n,p}^{(i)}$ from equations (47) and (49).
18:Compute denoised and damped channel variance $\hat{\sigma}_{h:{n,p}}^{2(i)}$ from equations (48) and (50).
19:Data Estimation: $\forall n,m$
20:Compute auxiliary variables $\tilde{\gamma}_{x:{m,n}}^{(i)}$ and $\hat{y}^{(i)}_{x_{m}:n,p}$ from eqs. (32) and (34), respectively.
21:Compute soft signal $\tilde{y}_{x:{m,n}}^{(i)}$ from equation (30).
22:Compute soft signal variance $\tilde{\sigma}^{2(i)}_{x:{m,n}}$ from equation (III-A1).
23:Compute extrinsic data belief $\tilde{x}_{n,m}^{(i)}$ and its variance $\tilde{\sigma}_{\tilde{x}:{n,m}}^{2(i)}$ from equation (39).
24:Compute denoised and damped data estimates $\hat{x}_{n,m}^{(i)}$ from equations (44) and (45).
25:Compute denoised and damped data estimate variances $\sigma_{x:{n,m}}^{2(i)}$ from equations (29a) and (46).
26:end for
Similarly to the above, in order to obtain the channel estimate belief, observe that the likelihood in equation (43) is a Gaussian-product distribution, with the mean of $p_{h}(h)$ equal to zero, which straightforwardly yields [39, 40]
$\hat{h}_{n,p}^{(i)}=\frac{\sigma^{2}_{h}\tilde{h}_{n,p}^{(i)}}{\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}+\sigma^{2}_{h}},$ (47)
and
$\hat{\sigma}^{2(i)}_{h:{n,p}}=\frac{\sigma^{2}_{h}\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}}{\tilde{\sigma}_{\tilde{h}:{n,p}}^{2(i)}+\sigma^{2}_{h}},$ (48)
which are respectively damped via
$\hat{h}_{n,p}^{(i)}=\beta_{h}\hat{h}_{n,p}^{(i)}+(1-\beta_{h})\hat{h}_{n,p}^{(i-1)}$ (49)
and
$\sigma_{h:{n,p}}^{2(i)}=\beta_{h}\sigma_{h:{n,p}}^{2(i)}+(1-\beta_{h})\sigma_{h:{n,p}}^{2(i-1)}.$ (50)
The proposed PBiGaBP-based JCDE algorithm is summarized as a pseudo code in Algorithm 1.
III-B Performance Analysis
[…]
First, we compare in Figure 4 the communications and channel estimation performances, in terms of BER and normalized mean square error (NMSE), respectively, achieved by PBiGaBP employed over OTFS, OFDM and AFDM waveforms. In this comparison, all systems employ an equal number $N_{p}$ of pilot symbols, although each system has its own structure (see Section II and Figure 2). The results show that PBiGaBP-based JCDE with AFDM and OTFS outperforms a system employing the same proposed technique but employing a legacy OFDM waveform, both in terms of BER and NMSE, with a slight advantage of AFDM over the OTFS alternative.
We highlight that a comparison between the proposed JCDE algorithm and other techniques from related literature is not offered in Figure 4, as we are unaware of any alternative JCDE solution for AFDM systems. Therefore, Figure 4 compares the proposed PBiGaBP-based JCDE against the GaBP alternative to the ML-based AFDM channel estimation method from [20].
We emphasize, however, that the technique in [20] is not a JCDE scheme, being rather strictly pilot-based, such that the latter SotA method enjoys, in this comparison, the advantage of not dealing with the additional burden of performing data detection, which the proposed algorithm does. In addition, as indicated in the legend of the figure, the comparison awards the SotA approach a power advantage, by means of allowing the latter to employ pilot signals with higher power than that used in the proposed method.
The results show that the proposed PBiGaBP-based JCDE scheme outperforms the GaBP-based method in [20] even when the latter employs a pilot with 10 times as higher power, which is a consequence of the fact that the proposed scheme takes advantage not only of the pilot, but also payload symbols to estimate the channel. The results also motivates us hereafter to no longer consider OFDM, as it is clear that this waveform is outperformed by OTFS and AFDM in doubly-dispersive channels.
(a) BER Performance.
(b) NMSE Performance.
Figure 3: Performance of OFDM, OTFS and AFDM at the UE via the proposed PBiGaBP algorithm compared to the SotA methods.
Figure 4: BER Performance of single pilot AFDM at the UE via the proposed PBiGaBP algorithm compared to the SotA methods.
IV Radar Parameter Estimation in Doubly Dispersive Channels
Having demonstrated the effectiveness of our proposed PBiGaBP-based JCDE scheme for AFDM, we next consider the radar parameter estimation problem [^14] For the sake of simplicity, we concentrate here on the parameter estimation – as opposed to target acquisition/detection – part of the radar problem, such that the total number of targets is assumed to be known [41], and the target association problem is assumed to be solvable [42]. It is easy to perceive from the approach introduced, however, that an extension of the work here presented to joint detection and parameter estimation, such as performed in [41], is possible. That will, however, be pursued in a follow-up work., also referred to as DD domain channel estimation problem [43]. To that end, a typical DD channel [30] is hereafter assumed, characterized by a maximum normalized delay spread $\text{max}(\ell_{p})$ and a corresponding digital normalized Doppler spread $\text{max}(f_{p})$, satisfying the relations $\text{max}(\ell_{p})<<N$ and $\text{max}(f_{p})<<N$.
Let us start by pointing out that the channel impulse response given in equation (1) can also be represented in the DD domain [13] as
$h(\tau,\nu)=\sum_{p=0}^{P}h_{p}\delta(\tau-\tau_{p})\delta(\nu-\nu_{p}),\vspace{-0.5ex}$ (51)
which for a finite number $P$ of paths, and under the aforementioned bounding assumption on delay and Doppler parameters, can be rewritten as
$h(\tau,\nu)=\sum_{k=0}^{K_{\tau}-1}\sum_{d=0}^{D_{\nu}-1}h_{k,d}\delta(\tau-\tau_{k})\delta(\nu-\nu_{d}),$ (52)
where $K_{\tau}$ and $D_{\nu}$ are numbers large enough to define a sufficiently fine grid discretizing the region determined by the maximum delay and Doppler values.
Notice that if the resolution of such a grid is sufficiently fine, the only non-zero terms of the sum in equation (52) are those where both $\tau_{k}\approx\tau_{p}$ and $\nu_{d}\approx\nu_{p}$, which in turn implies that estimating these radar parameters amounts to estimating the $P$ channel gains such that $h_{k,d}\neq 0$. This suggests that the radar parameter estimation problem can be formulated as a canonical sparse signal recovery problem [44] as follows.
First, rewriting the time-continuous model of equation (52) in terms of the sampled equivalent in matrix form yields
$\mathbf{\bar{H}}\triangleq\sum_{k=0}^{K_{\tau}-1}\sum_{d=0}^{D_{\nu}-1}h_{k,d}\cdot\mathbf{\Phi}_{k}\cdot\bm{\Omega}^{f_{d}}\cdot\mathbf{\Pi}^{\ell_{k}}\in\mathbb{C}^{N\times N},$ (53)
where we highlight that the model described by equation (53) is identical to the equivalent channel implicitly defined in equation (4), only with the summation rewritten over the DD grid, as opposed to the channel paths.
Then, the generalized receive signal model given in equation (27) can be rewritten as
$\mathbf{y}\!=\!\!\sum_{k=0}^{K_{\tau}\!-\!1}\sum_{d=0}^{D_{\nu}-1}\!\!\underbrace{\mathbf{\Gamma}_{k,d}\cdot\mathbf{x}}_{\mathbf{e}_{k,d}\in\mathbb{C}^{N\times 1}}\!\!\!\cdot h_{k,d}+\tilde{\mathbf{w}}=\mathbf{E}\cdot\mathbf{h}+\tilde{\mathbf{w}}\in\mathbb{C}^{N\times 1},\vspace{-1.5ex}$ (54a)
where we implicitly defined the vectors $\mathbf{e}_{k,d}\triangleq\mathbf{\Gamma}_{k,d}\cdot\mathbf{x}$, and explicitly identified the dictionary matrix $\mathbf{E}\in\mathbb{C}^{N\times K_{\tau}D_{\nu}}$, and sparse channel vector $\mathbf{h}\in\mathbb{C}^{K_{\tau}D_{\nu}\times 1}$, respectively defined as
$\mathbf{E}\!\triangleq\![\mathbf{e}_{0,0},\dots,\mathbf{e}_{0,D_{\nu}\!-\!1},\dots,\mathbf{e}_{K_{\tau}\!-\!1,0},\dots,\mathbf{e}_{K_{\tau}\!-\!1,D_{\nu}\!-\!1}],\vspace{-1ex}$ (54b)
$\mathbf{h}\!\triangleq\![h_{0,0},\dots,h_{0,D_{\nu}\!-\!1},\dots,h_{K_{\tau}\!-\!1,0},\dots,h_{K_{\tau}\!-\!1,D_{\nu}\!-\!1}]^{\mathsf{T}}\!\!\!.\vspace{-1ex}$ (54c)
In summary, the problem of estimating the set of radar parameters $\{\tau_{p},\nu_{p}\},\forall p$ reduces to estimating the sparse channel vector $\mathbf{h}$, with the delays and Doppler shifts obtained from the corresponding indices $(k,d)$ where $h_{k,d}\neq 0$, given the received signal $\mathbf{y}$ and the dictionary matrix $\mathbf{E}$. As a consequence of the approach, the assumption that $P$ is known is relaxed into the assumption that the channel paths are orthogonal in the delay-Doppler grid, such that $P$ can be unambiguously inferred from $\mathbf{h}$. Finally, the method also implies the assumption that the transmit vector $\mathbf{x}$ is known at the radar receiver, which in turn translates to a number of possible scenarios [13], including those described below.
Monostatic RPE: A first scenario is the classic monostatic case, in which the BS processes round-trip signals reflected by surrounding scatters (targets) in order to extract their distances and velocities. Here, a main practical challenge is the self-interference caused by signals transmitted by the BS at the same time when reflected signals return. The issue is often ignored in related literature on OFDM- and OTFS-based monostatic RPE [8, 18], implying a strong and not entirely practical assumption that self-interference is perfectly avoided by FD technologies. To emphasize this point, it has been shown that AFDM has an advantage over OFDM and OTFS as it allows for the concentration of a single pilot symbol within a transmission block [^15] Observe from equations (25) and (26) that unlike OFDM and OTFS, the AFDM receive signal occupies the entire time-frequency span even when a single entry in the transmit vector $\mathbf{x}$ is non-zero, such that concentrating piloting to a single symbol does not sacrifice the DoF of the estimator., as illustrated in Figure 2, with the corresponding interference mitigated via a de-chirping operation [20]. We will therefore focus on this case.
User-centric RPE: In this case, the BS first transmits a pilot symbol vector used by the UE to perform RPE, and subsequently transmits data from which the UE proceeds to perform JCDE as described in Section III in order to track the channel and decode the information. Under the assumption that delays and Dopplers change at a rate slower than $h_{p}$ [45], the alternation of JCDE and RPE stages would enable the system to minimize the overhead associated with frequent pilot transmissions. Due to the lack of space, this scenario will not be further developed here, but we have shown in [27] that a well-designed JCDE algorithm is capable of tracking time-varying channels even without periodic pilot transmissions, as long as the BER of the system is maintained sufficiently low, $e.g.$ via adequate channel coding. It is not difficult to envision that by combining the contribution of this article with the one in [27], the user-centric RPE scenario can be addressed.
Bistatic RPE: This refers to the case when multiple (passive) BSs perform RPE using signals transmitted by one (active) BS. The scenario can be similar either to the latter user-centric case, if the passive BSs have no knowledge of the transmit vector $\mathbf{x}$, or to the monostatic case, if $\mathbf{x}$ can be assumed known to all passive BSs. It is clear, therefore, that this scenario is a generalization of both, which is best addressed in a dedicated follow up article, especially since the latter assumption requires the BSs to be interconnected via a fronthaul network, which in turn might be subjected to impairments such as delays, noise, capacity limitations etc.
IV-A Reference SotA Method: SBL via Expectation Maximization
Before we introduce our contribution, let us briefly review an effective SotA channel estimation approach, which is here adapted to solve the RPE problem, and which will be used as a reference for performance assessment. The technique, which is based on the SBL framework, was proposed in [30] for OTFS systems, but can also be employed to AFDM thanks to the unified model of Section II and the resulting sparse linear representation of equation (54). To see this, recall that the SBL approach models the sparse channel vector $\mathbf{h}$ as a parameterized Gaussian distribution of the form
$f(\mathbf{h};\mathbf{\Xi})=\prod_{m=0}^{K_{\tau}D_{\nu}-1}\frac{1}{\pi\xi_{m}}\text{exp}\bigg{(}-\frac{|h_{m}|^{2}}{\xi_{m}}\bigg{)},$ (55)
where $\xi_{m}$ denotes the unknown hyperparameter corresponding to the $m$-th component of the vector $\mathbf{h}$, collected into the matrix $\mathbf{\Xi}\triangleq\text{diag}\big{(}\{\xi_{m}\}_{m=0}^{K_{\tau}D_{\nu}-1}\big{)}\in\mathbb{R}_{+}^{K_{\tau}D_{\nu}\times K_{\tau}D_{\nu}}$.
Then, estimates $\mathbf{\!\hat{\;h}}$ of the sparse vector and its corresponding hyperparameter matrix $\hat{\mathbf{\Xi}}$ are computed iteratively leveraging the EM method, which requires knowledge of the noise covariance matrix associated with the received signal model.
For the case of OTFS, from the definition of $\tilde{\mathbf{w}}_{\text{OTFS}}$ given just after equation (18) and under the assumption of rectangular pulse-shaping, we readily have
$\mathbf{R_{w}}\!\triangleq\!\mathbb{E}\Big{[}\tilde{\mathbf{w}}_{\text{OTFS}}\!\cdot\!\tilde{\mathbf{w}}_{\text{OTFS}}^{\mathsf{H}}\Big{]}\!\!=\!\mathbb{E}\Big{[}\big{[}(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})\mathbf{w}\big{]}\!\!\cdot\!\!\big{[}(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})\mathbf{w}\big{]}^{\mathsf{H}}\Big{]}$
$=(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})\!\cdot\!\sigma^{2}_{w}\mathbf{I}_{N}\!\cdot\!(\mathbf{F}_{M}\!\otimes\!\mathbf{I}_{K})^{\mathsf{H}}\!=\!\sigma^{2}_{w}\Big{[}\mathbf{F}_{M}\mathbf{F}_{M}^{\mathsf{H}}\!\otimes\!\mathbf{I}_{K}\mathbf{I}_{K}^{\mathsf{H}}\Big{]}$
$=\sigma^{2}_{w}\Big{[}\mathbf{I}_{M}\otimes\mathbf{I}_{K}\mathbf{I}_{K}^{\mathsf{H}}\Big{]}=\sigma^{2}_{w}\mathbf{I}_{N}\in\mathbb{C}^{N\times N}.$ (56)
In turn, taking into account the definition of $\tilde{\mathbf{w}}_{\text{AFDM}}$ given after equation (25), we have for the AFDM case
$\mathbf{R_{w}}\!\triangleq\!\mathbb{E}\Big{[}\tilde{\mathbf{w}}_{\text{AFDM}}\!\cdot\!\tilde{\mathbf{w}}_{\text{AFDM}}^{\mathsf{H}}\Big{]}\!\!=\!\mathbb{E}\Big{[}\!\big{[}(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)\mathbf{w}\big{]}\!\!\cdot\!\!\big{[}(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)\mathbf{w}\big{]}^{\mathsf{H}}\!\Big{]}$
$=\mathbb{E}\Big{[}\big{[}(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)\mathbf{w}\big{]}\!\cdot\!\big{[}\mathbf{w}^{\mathsf{H}}(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)^{\mathsf{H}}\big{]}\Big{]}$
$=(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)\!\cdot\!\sigma^{2}_{w}\mathbf{I}_{N}\!\cdot\!(\!\mathbf{\Lambda}_{2}\mathbf{F}_{\!N}\mathbf{\Lambda}_{1}\!)^{\mathsf{H}}\!=\!\sigma^{2}_{w}\mathbf{I}_{N}\in\mathbb{C}^{N\times N}.$
A pseudocode summary of the SBL-based RPE scheme is offered in Algorithm 2.
Algorithm 2 SBL-based Radar Parameter Estimation [30]
Input: Receive signal $\mathbf{y}$, dictionary matrix $\mathbf{E}$, noise covariance matrix $\mathbf{R_{w}}$, and stopping criteria parameters $\epsilon$ and $i_{\text{max}}$.
Output: Sparse channel estimate vector $\mathbf{\!\hat{\;h}}$.
Initialization: Set counter to $i=0$, and initial hyperparameters to $\hat{\mathbf{\Xi}}^{(0)}=\mathbf{I}_{K_{\tau}D_{\nu}}$ and $\hat{\mathbf{\Xi}}^{(-1)}=\mathbf{0},$
1:while $||\hat{\mathbf{\Xi}}^{(i)}-\hat{\mathbf{\Xi}}^{(i-1)}||^{2}_{F}>\epsilon$ and $i<i_{\text{max}}$ do
2:$i\leftarrow i+1$
3: E-step: Compute aposteriori covariance and estimate vector
4:            $\bm{\Sigma}^{(i)}=\big{[}\mathbf{E}^{\mathsf{H}}\mathbf{R}^{-1}_{\mathbf{w}}\mathbf{E}+(\hat{\mathbf{\Xi}}^{(i-1)})^{-1}\big{]}^{-1}$,
5:            $\mathbf{\!\hat{\;h}}^{\!(i)}=\bm{\Sigma}^{(i)}\mathbf{E}^{\mathsf{H}}\mathbf{R}_{\mathbf{w}}^{-1}\mathbf{y}$.
6: M-step: Update the hyperparameters $\hat{\mathbf{\Xi}}^{(i)}$, with
7:            $\hat{\xi}_{m}^{(i)}=\sigma_{m,m}^{(i)}+\big{|}\hat{h}_{m}^{(i)}\big{|}^{2}\!\!,\;\forall\,m$.
8:end while
IV-B Proposed Method: PDA-based Message Passing Estimator
In this section, we propose a novel PDA-based RPE scheme designed under the assumption that the prior distributions of the elements $\hat{h}_{m}$ of the sparse channel vector estimates $\mathbf{\!\hat{\;h}}$ are of Bernoulli-Gaussian type. In other words, we model the unknown channel to be estimated at the $i$-th iteration of the proposed scheme as
$\hat{h}_{m}^{(i)}\sim p_{\,\text{h}_{m}}(\text{h}_{m};\bm{\theta}^{(i)}),\vspace{-2ex}$ (58)
with
$\!\!p_{\,\text{h}_{m}}(\text{h}_{m};\bm{\theta}^{(i)})\triangleq\!(1\!-\!{\rho}^{(i)})\delta(h_{m})\!+\!{\rho}^{(i)}\mathcal{CN}\big{(}h_{m};{\bar{h}}^{(i)},{\bar{\sigma}}^{(i)}\big{)},\!\!\vspace{-0.5ex}$ (59)
where $\bm{\theta}^{(i)}\triangleq[{\rho}^{(i)},{\bar{h}}^{(i)},{\bar{\sigma}}^{(i)}]$ carries all three parameters of the distribution, namely, sparsity rate, mean and variance.
As implied by the notation, the parameter vector $\bm{\theta}^{(i)}$ must be updated iteratively, which in similarity to the SBL approach of [30] can be accomplished via an EM algorithm. Unlike the aforementioned SotA method, however, the estimates of $\mathbf{h}$ in the proposed scheme are obtained via a message-passing algorithm, which is described in the sequel.
For later convenience, we define the soft replica (i.e., tentative estimate) of $h_{m}$ as $\big{\{}\hat{h}_{m}^{(i)}\big{\}}$ such that its MSE can be expressed as
$\hat{\sigma}_{h:m}^{2(i)}\triangleq\mathbb{E}_{\text{h}_{m}}\big{[}|h_{m}-\hat{h}_{m}^{(i)}|^{2}\big{]}.$ (60)
Following steps similar to those in Section III-A, we have
IV-B1 Soft IC
The soft IC expression corresponding to an estimate of $h_{m}$, is given by
$\tilde{\mathbf{y}}_{h:m}^{(i)}\!=\mathbf{y}\!-\!\!\!\sum_{q\neq m}\!\!\mathbf{e}_{q}\hat{h}_{q}^{(i)}\!=\!\mathbf{e}_{m}h_{m}\!+\!\!\!\!\!\!\!\overbrace{\sum_{q\neq m}^{K_{\tau}D_{\nu}}(\mathbf{e}_{q}h_{q}\!-\!\mathbf{e}_{q}\hat{h}_{q}^{(i)})\!+\!\tilde{\mathbf{w}}}^{\text{residual interference+noise component}}\!\!\!,\!\!$ (61)
where $\mathbf{e}_{m}$ is the $m$-th column of the dictionary matrix $\mathbf{E}$.
It follows from the central limit theorem (CLT) that, under large-system conditions, the residual interference-plus-noise component can be approximated by a multivariate complex Gaussian variate. In other words, the vector Gaussian approximation (VGA) can be applied such that the conditional PDF of the beliefs $\tilde{\mathbf{y}}_{h:m}^{(i)}$, given $h_{m}$, can be expressed as
$\!\!\tilde{\mathbf{y}}_{h:m}^{(i)}\!\!\sim\!p_{\textbf{y}|\text{h}_{m}}\!(\textbf{y}|h_{m}\!)\!\propto\!\text{exp}\!\Big{[}\!-\!\big{(}\!\textbf{y}\!-\!\mathbf{e}_{m}h_{m}\!\big{)}^{\mathsf{H}}\mathbf{\Sigma}^{-1(i)}_{m}\!\big{(}\!\textbf{y}\!-\!\mathbf{e}_{m}h_{m}\!\big{)}\!\Big{]}\!,\!\!\!$ (62)
where y is an auxiliary variable, and the conditional covariance matrix $\mathbf{\Sigma}_{m}^{(i)}$ is given by
$\mathbf{\Sigma}_{m}^{(i)}$ $\triangleq\mathbb{E}_{\textbf{h,$\tilde{\mathbf{w}}$}|\hat{h}_{m}\neq h_{m}}\bigg{[}\big{(}\tilde{\mathbf{y}}_{h:m}^{(i)}-\mathbf{e}_{m}h_{m}\big{)}\big{(}\tilde{\mathbf{y}}_{h:m}^{(i)}-\mathbf{e}_{m}h_{m}\big{)}^{\mathsf{H}}\bigg{]}$
$=\sum_{q\neq m}^{K_{\tau}D_{\nu}}\hat{\sigma}_{h:q}^{2(i)}\mathbf{e}_{q}\mathbf{e}_{q}^{\mathsf{H}}+N_{0}\mathbf{I}_{N},$ (63)
with $N_{0}$ denoting the noise power.
IV-B2 Belief Generation
The beliefs associated with the estimate of the $m$-th channel entry $h_{m}$ can be obtained by combining the contributions of all soft IC beliefs $\tilde{\mathbf{y}}_{h:m}^{(i)}$, under the PDF
$p_{\text{h}|h_{m}}(\text{h}|h_{m})\propto\text{exp}\Big{[}-\tfrac{|\text{h}-\tilde{h}_{m}^{(i)}|^{2}}{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}}\Big{]},$ (64)
which yields
$\tilde{h}_{m}^{(i)}\triangleq\frac{1}{\eta_{m}^{(i)}}\mathbf{e}_{m}^{\mathsf{H}}\mathbf{\Sigma}^{-1(i)}\tilde{\mathbf{y}}_{h:m}^{(i)}\;\;\text{and}\;\;\tilde{\sigma}_{\tilde{h}:m}^{2(i)}\triangleq\frac{1-\eta_{m}^{(i)}\hat{\sigma}_{h:m}^{2(i)}}{\eta_{m}^{(i)}},$ (65a)
where $\eta_{m}^{(i)}$ is a normalization factor defined as
$\eta_{m}^{(i)}\triangleq\mathbf{e}_{m}^{\mathsf{H}}\mathbf{\Sigma}^{-1(i)}\mathbf{e}_{m},$ (65b)
and the common conditional covariance matrix [^16] The matrix inversion lemma [24] is used in the derivation of equation (65) and by consequence (65b), such that the same inverse matrix $\mathbf{\Sigma}^{(i)}$ can be used instead of $\mathbf{\Sigma}_{m}^{(i)}$. is given by
$\mathbf{\Sigma}^{(i)}\triangleq\sum_{m=1}^{K_{\tau}D_{\nu}}\hat{\sigma}_{h:m}^{2(i)}\mathbf{e}_{m}\mathbf{e}_{m}^{\mathsf{H}}+N_{0}\mathbf{I}_{N}.$ (65c)
IV-B3 Soft RG
Similarly to Section III, except that approximations hold under VGA as opposed to SGA, the soft replicas of $h_{m}$ can be inferred from the conditional expectation given the extrinsic beliefs and that the effective noise components in $\hat{h}_{m}^{(i)},\forall m$ are uncorrelated, which yields
$p_{h_{m}\!|\text{h}}(h_{m}|\text{h};\bm{\theta}^{(i)})\!=\!\frac{p_{\text{h}|h_{m}}\!(\text{h}|h_{m};\!\tilde{h}_{m}^{(i)}\!,\tilde{\sigma}_{\tilde{h}:m}^{2(i)})\;p_{h_{m}}\!(h_{m};\bm{\theta}^{(i)})}{\!\int_{h^{\prime}_{m}}\!p_{\text{h}|h_{m}}\!(\text{h}|h^{\prime}_{m};\!\tilde{h}_{m}^{(i)},\tilde{\sigma}_{\tilde{h}:m}^{2(i)})\;p_{h_{m}}\!(h^{\prime}_{m};\!\bm{\theta}^{(i)})}.$ (66)
Next, leveraging the assumption that $h_{m}$ follows a Bernoulli-Gaussian distribution, and using the Gaussian-PDF multiplication rule [40], equation (66) can be rewritten as [46]
$p_{h_{m}|\text{h}}(h_{m}|\text{h};\bm{\theta}^{(i)})\!=\!(1\!-\!\hat{\rho}_{m}^{(i)})\delta(h_{m})\!+\!\hat{\rho}_{m}^{(i)}\mathcal{CN}(h_{m};\hat{h}_{m}^{(i)},\hat{\sigma}_{h:m}^{2(i)}),$ (67)
where
$\vspace{-1ex}\hat{\rho}_{m}^{(i)}\triangleq\Bigg{(}\frac{1\!-\!{\rho}^{(i)}}{{\rho}^{(i)}}\frac{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}\!+\!{\bar{\sigma}}^{(i)}}{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}}\>e^{-\frac{|\tilde{h}_{m}^{(i)}|^{2}}{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}}+\frac{|\tilde{h}_{m}^{(i)}-{\bar{h}}^{(i)}|^{2}}{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}+{\bar{\sigma}}^{(i)}}}\!\!+1\Bigg{)}^{\!\!-1}\!\!\!\!,$ (68)
with
$\hat{h}_{m}^{(i)}\triangleq\frac{{\bar{\sigma}}^{(i)}\tilde{h}_{m}^{(i)}+\tilde{\sigma}_{\tilde{h}:m}^{2(i)}{\bar{h}}^{(i)}}{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}+{\bar{\sigma}}^{(i)}}\;\;\text{and}\;\;\hat{\sigma}_{h:m}^{2(i)}\triangleq\frac{{\bar{\sigma}}^{(i)}\tilde{\sigma}_{\tilde{h}:m}^{2(i)}}{\tilde{\sigma}_{\tilde{h}:m}^{2(i)}+{\bar{\sigma}}^{(i)}}.$ (69)
Finally, adhering to the prior defined for the channel coefficient $h_{p}$, we choose wlg, a denoiser without the mean parameter, namely, ${\bar{h}}^{(i)}=0$ in equation (59) and consequently, in equation (69) [^17] For completeness, the EM parameter update detailed in the consequent section will be derived for all the parameters.. From (67), the soft replica $\hat{h}_{m}^{(i)}$ and its MSE $\hat{\sigma}_{h:m}^{2(i)}$ can be in general obtained from the conditional expectation as [^18] Note the already incorporated damping procedure as also done in Section III to prevent convergence to local minima due to incorrect hard-decision replicas.
$\hat{h}_{m}^{(i)}=\tilde{\beta}_{h}\hat{\rho}_{m}^{(i)}\hat{h}_{m}^{(i)}+(1-\tilde{\beta}_{h})\hat{h}_{m}^{(i-1)},$ (70a)
$\hat{\sigma}_{h:m}^{2(i)}=\tilde{\beta}_{h}\big{[}(1-\hat{\rho}_{m}^{(i)})\hat{\rho}_{m}^{(i)}|\hat{h}_{m}^{(i)}|^{2}+\hat{\rho}_{m}^{(i)}\hat{\sigma}_{h:m}^{2(i)}\big{]}+(1-\tilde{\beta}_{h})\big{[}\hat{\sigma}_{h:m}^{2(i-1)}\big{]}.$ (70b)
IV-B4 Parameter Update via EM
In order to update the parameter set $\bm{\theta}$ of the Bernoulli-Gaussian distribution at each iteration, the EM algorithm is utilized. In particular, we employ EM as an iterative parameter tuning technique to obtain the parameter vector $\bm{\theta}$ that maximizes the likelihood function $p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})$, where $\mathbf{y}$ is defined in equation (54a), with $\bm{\theta}$ as given in equation (59).
To this end, we follow [47] and first convert $p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})$ to a tractable form by utilizing latent variables, marginalizing over an arbitrary PDF $\hat{p}_{\mathbf{{h}}}(\mathbf{h})$ and using the log-likelihood function $\text{ln}\big{[}p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})\big{]}$, reformulated so as to highlight the differential entropy and Kullback-Leibler (KL) divergence terms $\psi(\mathbf{h})$ and $\mathrm{D}_{\text{KL}}\big{(}\hat{p}_{\mathbf{{h}}}(\mathbf{h})||p_{\mathbf{h}|\mathbf{y},\bm{\theta}}(\mathbf{h}|\mathbf{y},\bm{\theta})\big{)}$, respectively, namely
$\text{ln}\big{[}p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})\big{]}\propto\int_{\mathbf{h}}\Big{[}\hat{p}_{\mathbf{{h}}}(\mathbf{h})\,\text{ln}\,p_{\mathbf{{h}},\mathbf{y},\bm{\theta}}(\mathbf{h},\mathbf{y},\bm{\theta})+\psi(\mathbf{h})$ (71)
$+\mathrm{D}_{\text{KL}}\big{(}\hat{p}_{\mathbf{{h}}}(\mathbf{h})||p_{\mathbf{h}|\mathbf{y},\bm{\theta}}(\mathbf{h}|\mathbf{y},\bm{\theta})\big{)}\Big{]}.$
Next, let us define the auxiliary function
$\mathrm{J}\big{(}\hat{p}_{\mathbf{{h}}}(\mathbf{h}),\bm{\theta}\big{)}\triangleq\int_{\mathbf{h}}\Big{[}\hat{p}_{\mathbf{{h}}}(\mathbf{h})\,\text{ln}\,p_{\mathbf{{h}},\mathbf{y},\bm{\theta}}(\mathbf{h},\mathbf{y},\bm{\theta})+\psi(\mathbf{h})\Big{]}.$ (72)
Then, taking advantage of the non-negativity of the KL divergence, the following lower bound on the latter log-likelihood function, can be obtained as
$\text{ln}\big{[}p_{\mathbf{y}|\bm{\theta}}(\mathbf{y}|\bm{\theta})\big{]}\geq\mathrm{J}\big{(}\hat{p}_{\mathbf{{h}}}(\mathbf{h}),\bm{\theta}\big{)}.$ (73)
The EM algorithm summarized above allows us to maximize the log-likelihood function in a cost-effective manner by alternating between the E-step minimizing KL divergence in equation (71) and the M-step maximizing the lower bound of the log-likelihood in equation (73), given by
[…]
Regarding the E-step, notice that solving equation (74a) is equivalent to minimizing of $\hat{p}_{\mathbf{{h}}}(\mathbf{h})$ given a fixed $\bm{\theta}$, but equation (67) already computes the solution to this problem. In other words, the Bernoulli-Gaussian distribution in equation (67) is the closed-form solution of problem (74a).
In turn, the focus of the M-step is to maximize $\bm{\theta}$ for a given distribution $\hat{p}_{\mathbf{{h}}}(\mathbf{h})$. Using (67), the maximization problem in (74b) can be reformulated as
Algorithm 3 PDA-based Radar Parameter Estimation (Prop.)
Input: Receive signal $\mathbf{y}$, dictionary matrix $\mathbf{E}$, noise power $N_{0}$, number of paths $P$, average channel power per path $\sigma_{h}^{2}$, maximum number of iterations $i_{\text{max}}$ and damping factor $\tilde{\beta}_{h}$.
Output: Estimates $\hat{\tau}_{p}$ and $\hat{\nu}_{p}$ extracted from the non-zero indices of the sparse channel estimate vector $\!\hat{\;\mathbf{h}}$.
Initialization: Set counter to $i=0$, set initial distribution parameters ${\rho}^{(0)}=P/(K_{\tau}D_{\nu})$ and ${\bar{\sigma}}^{(0)}=1/P$, set average channel power per path $\sigma_{h}^{2}=1/(K_{\tau}D_{\nu})$, and set initial estimates $\hat{h}_{m}^{(0)}=0$ and ${\hat{\sigma}}_{h:m}^{(0)}=\sigma_{h}^{2},\forall m$.
1:for $i=1$ to $i_{\text{max}}$ do $\forall m$
2:Compute soft signal vectors $\tilde{\mathbf{y}}_{h:m}^{(i)}$ from equation (61).
3:Compute soft extrinsic channel beliefs $\tilde{h}_{m}^{(i)}$ and their variances $\tilde{\sigma}_{\tilde{h}:m}^{2(i)}$ from equation (65).
4:Compute denoised sparsity rates $\hat{\rho}_{m}^{(i)}$ from eq. (68).
5:Compute denoised channel estimates $\hat{h}_{m}^{(i)}$ and their variances $\hat{\sigma}_{h:m}^{2(i)}$ from equation (69).
6:Compute damped channel estimates $\hat{h}_{m}^{(i)}$ and variances $\hat{\sigma}_{h:m}^{2(i)}$ from equation (70).
7:Update distribution parameters ${\rho}^{(i)}$ and ${\bar{\sigma}}^{(i)}$ from eq. (76c).
8:end for
Compute the estimates $\hat{\tau}_{p}$ and $\hat{\nu}_{p}$ corresponding to the indices $m$ of the non-zero entries of $\!\hat{\;\mathbf{h}}$ in accordance to equation (54c).
$\bm{\theta}=\underset{\bm{\theta}^{\prime}}{\mathrm{arg\;max}}\;\mathbb{E}_{\mathbf{h}|\text{h}}\Big{\{}\text{ln}\big{[}p_{\mathbf{{h}},\mathbf{y},\bm{\theta}}(\mathbf{h},\mathbf{y},\bm{\theta}^{\prime})\big{]}|\;\text{h};\bm{\theta}^{(i)}\Big{\}},$ (75)
and solved efficiently via a Lagrange method, yielding the update rules
${\rho}^{(i)}=\frac{1}{K_{\tau}D_{\nu}}\sum_{m=1}^{K_{\tau}D_{\nu}}\hat{\rho}_{m}^{(i)},\vspace{-0.75ex}$ (76a)
${\bar{h}}^{(i)}=\frac{1}{{K_{\tau}D_{\nu}}\cdot{\rho}^{(i)}}\sum_{m=1}^{K_{\tau}D_{\nu}}\hat{\rho}_{m}^{(i)}\hat{h}_{m}^{(i)},\vspace{-0.75ex}$ (76b)
${\bar{\sigma}}^{(i)}=\frac{1}{{K_{\tau}D_{\nu}}\cdot{\rho}^{(i)}}\sum_{m=1}^{K_{\tau}D_{\nu}}\hat{\rho}_{m}^{(i)}\bigg{(}\big{|}\hat{h}_{m}^{(i)}-{\bar{h}}^{(i)}\big{|}^{2}+\hat{\sigma}_{h:m}^{2(i)}\bigg{)}.\vspace{-0.5ex}$ (76c)
A complete and compact description of the proposed PDA-based RPE scheme is given in the form of a pseudocode in Algorithm (3). Upon convergence of the above procedure, the estimate of the sparse vector $\mathbf{h}$ is known. Finally, the $(i,j)$ indices of $\mathbf{h}$ can be used to obtain the corresponding delays and Dopplers, and hence the corresponding ranges and velocities of the objects in the vicinity.
IV-C Performance Assessment
In order to evaluate the performance of the proposed PDA-based RPE algorithm, we consider the monostatic scenario depicted in Fig. 1. Under such a scenario, and in light of the results of Figure 4 – as well as related literature, $e.g.$[8] – which demonstrate that OFDM is outperformed by OTFS and AFDM in doubly-dispersive channels, we consider hereafter only OTFS and AFDM systems.
Our first set of results, shown in 20, are obtained for a single LoS reflected signal, from a single target, positioned at $15$ m and traveling at a velocity of $151$ km/h towards the BS, with the remark that since both the SotA SBL and the proposed PDA RPE algorithms are based on the sparse model described by the set of equations (54), the estimation algorithm themselves have no hard restrictions on the number of paths or targets to be estimated, as long as the grid modeled by the dictionary matrix $\mathbf{E}$ in equation (54b) is fine enough to ensure a sufficient sparsity of the vector $\mathbf{y}$ in equation (54a).
With that clarified, the motivation of assessing the performance of these algorithms under a single path/target case is to grant an advantage to the OTFS approach, where a larger number of targets only increases the weight of the assumption of no self-interference (not required in the AFDM case).
[…]
The performance metric adopted is the normalized root mean square error (RMSE), averaged over all targets and realizations, which is defined as
$\Upsilon\triangleq\mathbb{E}\Big{[}\frac{1}{P+1}\sum_{p=0}^{P}\frac{||\hat{\vartheta}_{p}-\vartheta_{p}||^{2}_{2}}{|\vartheta_{p}|}\Big{]},\vspace{-0.5ex}$ (77)
where $\vartheta_{p}$ denotes a given radar parameter (range or velocity), $\hat{\vartheta}_{p}$ its estimate, and the expectation is taken over the number of realizations of the algorithm.
Finally, Figure 20 also includes plots of the resolution limit [12] and the MF bound [48], obtained by executing Algorithm 3 for a single iteration, initialized with the true radar parameter values. The results clearly indicate that the proposed PDA method outperforms the SotA SBL approach both for range and velocity estimation, both under OTFS and AFDM waveforms, with the proposed method reaching the MF bound in the case of range estimates. It is also found, moreover, that AFDM yields slightly better performance than OTFS, and that all algorithms reach the resolution bound under a sufficiently large signal-to-noise ratio (SNR), indicating that both methods converge absolutely.
(a) Range Estimation Performance (FD system, single target).
(b) Velocity Estimation Performance (FD system, single target).
Figure 5: RMSE versus SNR  [^20] While many formulations for the SNR exist in literature, including radar-centric metrics as used in [8, 12], we use the typical definition hinging on $h_{p}$ as used in [20].performance of SotA SBL and proposed PDA monostatic RPE schemes over OTFS and AFDM waveforms in a FD (interference-free) scenario with a single LoS signal from a single target.
Figure 6: Convergence of SotA SBL and proposed PDA monostatic RPE schemes over OTFS and AFDM.
(a) Range Estimation Performance (SIC system, multiple targets).
(b) Velocity Estimation Performance (SIC system, multiple targets).
Figure 7: RMSE versus SNR performance of SotA SBL and proposed PDA monostatic RPE schemes over AFDM waveform, subjected to self-interference removed via de-chirping as in [20], in a scenario with multiple targets.
In order to clarify that convergence is also not an issue at low SNR, we offer in Figure 7 plots of the range and velocity normalized RMSE. It can be seen from those results that the proposed PDA-based method is advantageous compared to the SotA SBL approach also in terms of convergence speed.
Having established the overall advantage of AFDM over OTFS, and of the proposed PDA-based RPE algorithm over the SBL-based SotA alternative, we finally confront the two critical issues left unaddressed in the performance assessment so far, namely, the impact of self-interference and a larger number of targets to be estimated. To that end, we offer in Figure 7 plots equivalent to those shown in Figure 20, but this time exclusively for AFDM systems, and in scenarios with multiple targets and subjected to self-interference, mitigated via the technique proposed in [20].
V Conclusion
We offered two new ISAC-enabling contributions for the emerging AFDM waveform, namely, a PBiGaBP-based JCDE algorithm, which enables channel estimation with low pilot overhead, and a PDA-based RPE algorithm, which enables the BS to act as a monostatic radar, estimating the delay and Dopper parameters of multiple paths simultaneously. Using a generalized channel model under which OFDM, OTFS and AFDM can be directly compared, it was shown that the proposed algorithms outperform SotA alternatives and that AFDM is advantageous over OTFS and OFDM both for communications and sensing functionalities in doubly-dispersive channels.
Appendices/Illustrations
Figure 8: Diagram of Proposed PBiGaBP-based JCDE in Algorithm 1.
Figure 9: Diagram of Proposed PDA-based RPE Algorithm 3.